OpenStudy (anonymous):

write the equation in standard form of the line that contains the points(5,-2) and is a) parallel to the line 3x+4y=7 and b) perpendicular to the line 3x+4y=7

OpenStudy (gorv):

parallel line have same slope equation of line in standard form y=mx+c 3x+4y=7 4y=-3x+7 y=-3x/4+7

OpenStudy (gorv):

sorry that will be y=-3x/4+7/4

OpenStudy (anonymous):

really?